OpenStudy (anonymous):

Find the limit if it exists. lim x-> 0 csc3x/cotx

OpenStudy (anonymous):

did you rewrite using sines and cosines?

OpenStudy (anonymous):

should get \[\frac{\sin(x)}{\sin(3x)\cos(x)}\] and you don't have to think about the cosine, because \(\cos(0)=1\) so it is a matter of finding \[\lim_{x\to 0}\frac{\sin(x)}{\sin(3x)}\]